Circular No. 7791

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002B IN ANONYMOUS GALAXY C. Stubbs, University of Washington, reports, on behalf of the Super-

Macho microlensing survey team (including also A. Rest, R. Covarrubias, A. Becker, C. Smith, K. Olsen, N. Suntzeff, R. Hiriart, D. Welch, D. Lepischak, A. Clocchiati, B. Schmidt, and K. Cook), the detection with the Cerro Tololo 4-m telescope (+ MOSAIC II imager) on Jan. 7.20 UT of an apparent supernova (R ∼ 20.5) in a galaxy behind the Large Magellanic Cloud. The new object was not apparent in images taken on 2001 Nov. 22 (limiting mag R ∼ 23) and, based on the accumulated light curve with data points from multiple nights, appears to reach a peak magnitude of R ∼ 20.5 on 2002 Jan. 9.2. SN 2002B is located at α = 5^{h}40^{m}46^{s}.06, δ = −71^{o}51′15′′.1 (equinox 2000.0), which is ∼ 2′′.2 west of the host galaxy’s nucleus.

IM NORMAE T. Kato, Kyoto University, notes that the proposed identification of

IM Nor with 2U 1536–52 (cf. IAUC 7789) is incorrect, with QV Nor = 2U 1536–52 = 4U 1538–52 shown to be valid (cf. IAUC 3184, 3201; Kholopov 1985, General Catalogue of Variable Stars 2, 348). H. Yamaoka, Kyushu University, having looked at various Digitized Sky Survey images of the field, reports that IM Nor may be identical with the easternmost star (having red mag ∼ 17.0, via comparison with nearby USNO-A2.0 stars on a red plate, and blue mag ∼ 18.0 from a blue plate taken on 1975 July 7) of three aligned stars. Yamaoka adds that the star appears less bright on a Palomar blue plate taken on 1987 Apr. 25, noting that the object may be variable in quiescence. W. Liller, Viña del Mar, Chile, reports that this variable had brightened to broadband V = 7.84 on Jan. 14.3285 UT.

PECULIAR VARIABLE IN MONOCEROS C. Bailyn, Yale University, reports on YALO-consortium observations

of this variable (IAUC 7786): “Images with the ANDICAM instrument on the Yale 1-m telescope at Cerro Tololo taken on Jan. 11.139 UT reveal a bright infrared source with J, H, and K magnitudes < 7. Preliminary reductions suggest K = 5.9, H = 6.2, J = 6.6, but these values may be imprecise due to non-linearity effects at these high count rates.”

Visual magnitude estimates: Jan. 11.01 UT, 10.6 (K. Sárneczky, Piszkéstető, Hungary); 11.616, 10.5 (A. Pearce, Nedlands, Western Australia); 13.375, 10.4 (J. Bedient, Honolulu, HI); 14.361, 10.3 (Bedient).

2002 January 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green